OpenStudy (anonymous):

convert 141km to miles

OpenStudy (anonymous):

its 87.613 miles

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

convert 5 kg to pound

OpenStudy (anonymous):

11.023

OpenStudy (anonymous):

convert 6 kg to pound

OpenStudy (anonymous):

13.227

OpenStudy (anonymous):

convert 7 kg to pounds

OpenStudy (anonymous):

15.432

OpenStudy (anonymous):

convert 19 km to miles

OpenStudy (anonymous):

11.806 miles

OpenStudy (unklerhaukus):

\[1.6[\text{km}] \approx[\text{mi}]\\1\approx\frac{[\text{mi}]}{1.6[\text{km}] }\] \[141[\text{km}] =141[\text{km}]\times1\\\qquad\quad=141\color{red}{\cancel{\color{black}{[\text{km}]}}} \times\frac{[\text{mi}]}{1.6\color{red}{\cancel{\color{black}{[\text{km}]}}} }\\\qquad\quad =\frac{141}{1.6}[\text{mi}]\]